OpenStudy (anonymous):

If f(x) = x2 - 1, find f(3).

OpenStudy (anonymous):

8

OpenStudy (anonymous):

I'm guessing by x2 you meant x squared

OpenStudy (anonymous):

f(3)=3*2-1 y=6-1 y=5

OpenStudy (anonymous):

or 8 if is x^2

OpenStudy (anonymous):

i meant x squared

OpenStudy (anonymous):

(3)^2 - 1 = 9 - 1 = 8